OpenStudy (anonymous):

Help Please! y is directly proportional to the square of x. when y=5, x=4. find the value of y when x=8. (3marks)

OpenStudy (anonymous):

"Directly proportional" means y = k x^2. Substitute the first x, y combination to find k, then use the equation to find y for x = 8.

OpenStudy (anonymous):

why is k used? and can i find it by doing x^2 * y = k

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

"Directly proportional" means that when x doubles, so does y, but it does not tell you that y=x, but only that y = k x, where you have to find k. If Joe's weight is proportional to his waist size, we still don't know how to get his weight exactly from knowing his waist, until we learn k, the constant of proportionality. Here, y = k x^2 and can get k from k = y / x^2 once you have an x, y pair to plug into that equation.